# On compatibility of Koszul- and higher preprojective gradings

### v2

@@ 5.5. Preprojective gradings and cuts
We are now ready to state our main theorem for the $n$-representation infinite case.
@@
Let $R=R_{\bullet}$ be a graded basic Koszul algebra, so that $R_{1}$ is generated by nilpotent elements. If there exists a higher preprojective grading ${}_{\bullet}R$ on $R$, then there exists an automorphism that maps the higher preprojective grading to a higher preprojective cut. In particular, all higher preprojective gradings are, up to isomorphism, higher preprojective cuts.
@@
Combining 5.10 and 5.6, we obtain that the graded simples for the Koszul grading and the higher preprojective grading coincide. Therefore, we can apply 5.13 to see that $A={}_{0}R$ is Koszul. Then, by 2.22, there exists a Koszul grading on $\Pi_{n+1}(A)\simeq R$ so that the higher preprojective grading on $\Pi_{n+1}(A)$ is a cut. Finally, by 2.9, there exists an automorphism taking the new Koszul grading on $\Pi_{n+1}(A)\simeq R$ to the given grading $R_{\bullet}$. Hence, this automorphism maps ${}_{\bullet}R$ to a cut. ∎
@@
Let $R=R_{\bullet}$ be a Koszul algebra, so that $R_{1}$ is generated by nilpotent elements. If there exists a higher preprojective grading ${}_{\bullet}R$ on $R$ such that $R_{0}\subseteq{}_{0}R$, then there exists a higher preprojective cut realising $R_{\bullet}=\Pi_{n+1}(A)$ so that $A\simeq{}_{0}R$.
@@
By assumption, we can reduce $R_{\bullet}$ to a graded basic algebra $R_{\bullet}^{b}$ by some homogeneous idempotent $e\in R_{0}\subseteq{}_{0}R$. Since $e$ is preprojectively homogeneous, it follows that $eRe$ inherits a higher preprojective grading. By 5.14, we obtain a higher preprojective cut, and by 3.8 this gives the higher preprojective cut on $R_{\bullet}$. ∎
@@
Let $A$ be basic and $n$-representation infinite and let $B$ be an $n$-APR tilt of $A$. Then $A$ is Koszul if and only if $B$ is Koszul.
@@
If $A$ is Koszul, then so is $\Pi_{n+1}(A)$ by 2.17, and the higher preprojective grading is a cut for this grading. Hence, the graded simple modules for the two gradings coincide as ungraded modules. Transferring the grading to $\Pi_{n+1}(B)$, we get a Koszul grading on $\Pi_{n+1}(B)$ satisfying the assumptions of 5.13, showing that $B$ is Koszul. ∎
@@ Acknowledgements
We thank Steffen Oppermann for helpful discussions, and for suggesting the proof of a Lemma which was used in a previous version of this article. We also thank Martin Herschend for many helpful discussions. The second author is grateful to have been supported by Norwegian Research Council project 301375, “Applications of reduction techniques and computations in representation theory”.

### v3

@@ 5.5. Preprojective gradings and cuts
We are now ready to state our main theorem for the $n$-representation infinite case. In light of Section 5.3, we fix the following terminology.
Definition 5.15.
A higher preprojective grading $_{\bullet}R$ on some algebra $R$ is called acyclic if the finite-dimensional algebra $_{0}R$ has acyclic Gabriel quiver.
@@
Let $R=R_{\bullet}$ be a graded basic Koszul algebra, so that $R_{1}$ is generated by nilpotent elements. If there exists an acyclic higher preprojective grading $_{\bullet}R$ on $R$, then there exists an automorphism that maps this higher preprojective grading to a higher preprojective cut. In particular, all acyclic higher preprojective gradings are, up to isomorphism, higher preprojective cuts.
@@
Combining 5.11 and 5.6, we obtain that the graded simples for the Koszul grading and the higher preprojective grading coincide. Therefore, we can apply 5.14 to see that $A={}_{0}R$ is Koszul. Then, by 2.22, there exists a Koszul grading on $\Pi_{n+1}(A)\simeq R$ so that the higher preprojective grading on $\Pi_{n+1}(A)$ is a cut. Finally, by 2.9, there exists an automorphism taking the new Koszul grading on $\Pi_{n+1}(A)\simeq R$ to the given grading $R_{\bullet}$. Hence, this automorphism maps $_{\bullet}R$ to a cut. ∎
@@
Let $R=R_{\bullet}$ be a Koszul algebra, so that $R_{1}$ is generated by nilpotent elements. If there exists an acyclic higher preprojective grading $_{\bullet}R$ on $R$ such that $R_{0}\subseteq{}_{0}R$, then there exists a higher preprojective cut realising $R_{\bullet}=\Pi_{n+1}(A)$ so that $A\simeq{}_{0}R$.
@@
By assumption, we can reduce $R_{\bullet}$ to a graded basic algebra $R_{\bullet}^{b}$ by some homogeneous idempotent $e\in R_{0}\subseteq{}_{0}R$. Since $e$ is preprojectively homogeneous, it follows that eReeRe inherits a higher preprojective grading. By 5.16, we obtain a higher preprojective cut, and by 3.8 this gives the higher preprojective cut on $R_{\bullet}$. ∎
@@
Let $A$ be basic and $n$-representation infinite, and let $B$ be an $n$-APR tilt of $A$. Then $A$ is Koszul if and only if $B$ is Koszul.
@@
If $A$ is Koszul, then so is $\Pi_{n+1}(A)$ by 2.17, and the higher preprojective grading is a cut for this grading. Hence, the graded simple modules for the two gradings coincide as ungraded modules. Transferring the grading to $\Pi_{n+1}(B)$, we get a Koszul grading on $\Pi_{n+1}(B)$ satisfying the assumptions of 5.14, showing that $B$ is Koszul. ∎
@@ Acknowledgements
We thank Gustavo Jasso for pointing out an error in a previous version of this article. We thank Steffen Oppermann for helpful discussions, and for suggesting the proof of a Lemma which was used in a previous version of this article. We also thank Martin Herschend for many helpful discussions. The second author is grateful to have been supported by Norwegian Research Council project 301375, “Applications of reduction techniques and computations in representation theory”.